Solve the problem.
The sum of two numbers is 12. If one number is subtracted from the other, their difference is -2. Find the numbers.

Answer by sofiyac(983)

x+y=12
x-y=-2
***
x=12-y
12-y-y=-2
-2y=-17
y=7
x-7=-2
x=5
the numbers are 5 and 7